When Ruth emptied her piggy-bank, of nickels and dimes, she counted 84 coins in total. The value of the coins was $7.15. how many dimes did she have?

Answer by Alan3354(69443)

Same problem, different numbers.
===============================================================
Mary has 30 nickels and dimes totaling $2.40. How many of each does she have
============
n + d = 30 --------- # of coins
5n + 10d = 240 --------Value of coins
--> n + 2d = 48 ---------- 2nd eqn divided by 5
=============================
n + d = 30 --------- # of coins
n + 2d = 48
------------ Subtract
-d = -18
--> 18 dimes
12 nickels